# Negation-Aware Test-Time Adaptation for Vision-Language Models

### v1

@@ V-D1 Impact of Each Component
To study the influence of specific components in our method, we first carry out ablation studies on the MS-COCO Retrieval-Neg data with different training objects. Table VIII reports the performance of the Retrieval-neg and the MCQ-Neg tasks. From the results, we observe the following conclusions: 1) $\mathcal{L}_{ent}$ boosts performance through the negation-separated prediction, i.e., improving the R@1 of the base CLIP by 2.76% and 4.24% on text and image retrieval, respectively. However, naive entropy minimization fails to generalize the model to more difficult negation understanding scenarios, e.g., performance drops significantly on the MCQ task. 2) $\mathcal{L}_{sr}$ is crucial for learning generalizable negation understanding capabilities, increasing the MCQ accuracy of the $\mathcal{L}_{ent}$ variant from 28.69% to 45.91%. 3) NEAT achieves overall optimal performance when all the loss terms are employed, showing that all three components make contributions. In addition, we conduct experiments to verify the effectiveness of the proposed negation-separated prediction in Section IV-B. We compared with the KNN-based prediction [23] and report the zero-shot top-1 prediction accuracy. As shown in Fig. 5a, our refinement strategy substantially outperforms the KNN variant, which signifies that the simple feature-driven prediction is insufficient for handling negation.
@@ V-D2 Parameter Analysis
We next investigate the effect of the parameter $\lambda$ by plotting the online Recall@1 scores with incremental values on MSR-VTT. As shown in Fig. 5b, we observe that: 1) when using smaller $\lambda$ values, e.g., 111 and 222, the second term in Eq.(8) would dominate the optimization objective, leading to sub-optimal model performance. 2) Our method can achieve stable performance in a relatively larger range, i.e. 5205$\sim$ 205 ∼ 20, and shows no substantial performance variation even when using extremely large $\lambda$ values.
@@ V-E1 Embedding Similarity Analysis
To visually investigate the performance of our NEAT against negation, we illustrate pairwise similarity distributions of image-text embeddings from different VLMs. Specifically, we compared our NEAT (adapted on unlabeled MS-COCO Retrieval-Neg) with CLIP, ConCLIP, NegFull-finetuned CLIP, and TCR (adapted on unlabeled MS-COCO Retrieval-Neg) in Fig. 6. For each image in the test set of CIFAR10, we contrast it with the normal text, i.e., “a photo of the [CLASS]”, the negation-conditioned text, i.e., “a photo of the [CLASS] but not of the [CLASS]”, and the reversed negation-conditioned text, i.e., “a photo of the [CLASS] but not of the [CLASS]”, where [CLASS] is the true label and [CLASS] is a false one. As shown in Figs. 6a-6e, all VLMs could separate the positive and negative pairs apart enough. Although some methods are designed for negation understanding, e.g., ConCLIP and NEAT, their ability to handle affirmative statements remains intact. When tasked with negation-conditioned understanding, Figs. 6f-6j demonstrate that some VLMs exhibit significant overlap between positive and negative distributions, eg, CLIP and ConCLIP. In contrast, our NEAT could discriminate the true and false negative pairs successfully. While for the more challenging reversed negation-conditioned text, Figs. 6k-6o show that CLIP and TCR wrongly provide higher similarity scores for positive (the most irrelevant) pairs. Although some negation-enhanced models, e.g., ConCLIP and NegFull-finetuned CLIP, reduce similarity scores for positive pairs, they still maintain distributions comparable to those of negative pairs. Differently, our NEAT is the only approach that properly produces lower similarity scores for positives than negatives, which is consistent with the findings in Table V.
@@ VI Conclusion
In this paper, we study enhancing the negation understanding ability of VLMs through test-time adaptation. The key idea is to address the dual-concept shifts problem between affirmation and negation distributions. Specifically, we propose NEAT, a novel method that reduces distribution shift in consistent semantics while eliminating false distributional consistency in unrelated semantics. By adjusting only the lightweight normalization layers, our method could efficiently adapt VLMs to negation contexts during inference. Extensive experiments across multiple benchmarks spanning images, videos, and medical scenarios verify the effectiveness and generalization of our method.

### v2

@@ V-D1 Impact of Each Component
To study the influence of specific components in our method, we first carry out ablation studies on the MS-COCO Retrieval-Neg data with different training objects. Table IX reports the performance of the Retrieval-neg and the MCQ-Neg tasks. From the results, we observe the following conclusions: 1) $\mathcal{L}_{ent}$ boosts performance through the negation-separated prediction, i.e., improving the R@1 of the base CLIP by 2.76% and 4.24% on text and image retrieval, respectively. However, naive entropy minimization fails to generalize the model to more difficult negation understanding scenarios, e.g., performance drops significantly on the MCQ task. 2) $\mathcal{L}_{sr}$ is crucial for learning generalizable negation understanding capabilities, increasing the MCQ accuracy of the $\mathcal{L}_{ent}$ variant from 28.69% to 45.91%. 3) NEAT achieves overall optimal performance when all the loss terms are employed, showing that all three components make contributions. In addition, we conduct experiments to verify the effectiveness of the proposed negation-separated prediction in Section IV-B. We compared with the KNN-based prediction [25] and report the zero-shot top-1 prediction accuracy. As shown in Fig. 5a, our refinement strategy substantially outperforms the KNN variant, which signifies that the simple feature-driven prediction is insufficient for handling negation.
@@
TABLE IX: Ablation study of the proposed training objects.
$\mathcal{L}_{ent}$ $\mathcal{L}_{sr}$ $\mathcal{L}_{tri}$ I2T R@1 T2I R@1 MCQ
45.30 24.96 39.25
48.06 29.20 28.69
7.12 4.67 42.78
49.74 28.95 27.26
44.96 30.05 45.91
48.98 30.00 48.41
@@ V-D2 Parameter Analysis
We next investigate the effect of the parameter $\lambda$ by plotting the online Recall@1 scores with incremental values on MSR-VTT. As shown in Fig. 5b, we observe that: 1) when using smaller $\lambda$ values, e.g., $1$ and $2$, the second term in Eq.(8) would dominate the optimization objective, leading to sub-optimal model performance. 2) Our method can achieve stable performance in a relatively larger range, i.e. $5\sim 20$, and shows no substantial performance variation even when using extremely large $\lambda$ values.
@@ V-E1 Embedding Similarity Analysis
To visually investigate the performance of our NEAT against negation, we illustrate pairwise similarity distributions of image-text embeddings from different VLMs. Specifically, we compared our NEAT (adapted on unlabeled MS-COCO Retrieval-Neg) with CLIP, ConCLIP, NegFull-finetuned CLIP, and TCR (adapted on unlabeled MS-COCO Retrieval-Neg) in Fig. 6. For each image in the test set of CIFAR10, we contrast it with the normal text, i.e., “a photo of the [CLASS]”, the negation-conditioned text, i.e., “a photo of the [CLASS] but not of the [CLASS]”, and the reversed negation-conditioned text, i.e., “a photo of the [CLASS] but not of the [CLASS]”, where [CLASS] is the true label and [CLASS] is a false one. As shown in Figs. 6a-6e, all VLMs could separate the positive and negative pairs apart enough. Although some methods are designed for negation understanding, e.g., ConCLIP and NEAT, their ability to handle affirmative statements remains intact. When tasked with negation-conditioned understanding, Figs. 6f-6j demonstrate that some VLMs exhibit significant overlap between positive and negative distributions, eg, CLIP and ConCLIP. In contrast, our NEAT could discriminate the true and false negative pairs successfully. While for the more challenging reversed negation-conditioned text, Figs. 6k-6o show that CLIP and TCR wrongly provide higher similarity scores for positive (the most irrelevant) pairs. Although some negation-enhanced models, e.g., ConCLIP and NegFull-finetuned CLIP, reduce similarity scores for positive pairs, they still maintain distributions comparable to those of negative pairs. Differently, our NEAT is the only approach that properly produces lower similarity scores for positives than negatives, which is consistent with the findings in Table VI.
@@ VI Conclusion
In this paper, we study enhancing the negation understanding ability of VLMs through test-time adaptation. The key idea is to address the dual-concept shifts problem between affirmation and negation distributions. Specifically, we propose NEAT, a novel method that reduces distribution shift in consistent semantics while eliminating false distributional consistency in unrelated semantics. By adjusting only the lightweight normalization layers, our method could efficiently adapt VLMs to negation contexts during inference. Extensive experiments across multiple benchmarks spanning images, videos, and medical scenarios verify the effectiveness and generalization of our method. In future work, we plan to explore more negation understanding scenarios, e.g., remote sensing image retrieval with negation-conditioned queries, and tasks, e.g., visual question answering, broadening NEAT to handle these corresponding challenges.